OpenStudy (anonymous):

i have a math question.. what is 1+1+1+1+1+1+1+1+1+1+1+-2.5467833 ANSWER???????????????? ooh another hard one. 2+2

OpenStudy (anonymous):

8.4532167 and 2

OpenStudy (sissyedgar):

LOL... 13.5467833 4

OpenStudy (sissyedgar):

UH WHERE U GET THT ANSWER FROM DEADPOOL?

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

hmmm......... I used a calculator....

OpenStudy (sissyedgar):

I DID 2... BUT I DIDNT GET THT

OpenStudy (anonymous):

I got an answer in less than a second. Use google to search it.

OpenStudy (ilovewolf):

1+1+1+1+1+1+1+1+1+1+1=11 11+2.5467833= 13.5467833 and 2+2=4 >.<

OpenStudy (sissyedgar):

SEE!

OpenStudy (ilovewolf):

lol

OpenStudy (anonymous):

you really shouldnt need a calculator for 2+2....

OpenStudy (sissyedgar):

IK

OpenStudy (anonymous):

meh. o__o

OpenStudy (anonymous):

2+2 is fish duh

OpenStudy (anonymous):

ppphhh! LOL